OpenStudy (anonymous):

What is c if y=-3x^2+6x+c if x=1 and y=0 5,-2,-3,or 2

OpenStudy (anonymous):

you just need to substitute the values and then evaluate.

OpenStudy (anonymous):

y=-3x^2+6x+c (0)= -3*(1)^2+6(1) +c 0 = -3+6+c 0=3+c -3=c

OpenStudy (anonymous):

y=0 c= -3 y=5 c= -45 y=-2 c=-24 y=-3 c=-45 y=2 c=0

OpenStudy (anonymous):

@ Rockstar it is -3x sqaured

OpenStudy (anonymous):

yes i kno

OpenStudy (anonymous):

3x^2 = 3*x*x